what is the pressure exerted by a 50kg girl as she places her weight on the hell of one shoe if the heel area is 0.002m²


1
Expert's answer
2021-03-10T17:16:31-0500
"P=\dfrac{F}{A}=\dfrac{mg}{A},""P=\dfrac{50\ kg\cdot9.8\ \dfrac{m}{s^2}}{0.002\ m^2}=2.45\cdot10^5\ Pa."